A pool has dimensions of 13 ft long, 10 ft wide and 6 ft deep.
Perimeter around the pool?
Area of the bottom of the Pool?

Answer by TimothyLamb(4379)

---
perimeter = 2(13 + 10)
perimeter = 2(23)
perimeter = 46 ft
---
area = 13 * 10
area = 130 sq.ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php